OpenStudy (anonymous):

A circle is described by the equation (x-1/2)^2+(y-3/2)^2=4/9. What are the cordinates for the center of the circle and the length of the radius? A)(-1/2,-3/2), 2/3units B) (-1/2,-3/2), 4/9units C) (1/2,3/2), 4/9units D) (1/2,3/2), 2/3units

OpenStudy (anonymous):

plz explain

OpenStudy (mrnood):

you have shown in the last two questions we tackled that you already know how to do this. the formula is (x-a)^2+(y-b)2 = r^2 (a,b) is the centre. Just look at the equation you have. It is easy to get a and b from the 'brackets' r is simple sqrt(right hand side).... You CAN do this

OpenStudy (anonymous):

so is it b?

OpenStudy (mrnood):

and sqrt(m/n) = sqrt(m)/sqrt(n) which makes it very easy to get the value for the radius....

OpenStudy (mrnood):

PLEASE do not guess. You can do this. I do not confirm or deny whether your guess is correct.